# Investigating the Nested Structure of the Outflow from the Low Luminosity Protostar IRAS 16253-2429 using JWST and ALMA

### v1

@@ I Introduction
The process of star formation is governed by the complex interplay between infall, accretion, and feedback mechanisms, which determine the final mass of the star and the initial conditions for planet formation (Dunham et al., 2014; Federrath et al., 2014; Guszejnov et al., 2018, 2022; Fischer et al., 2023; Hsieh et al., 2023; Tobin & Sheehan, 2024; Hennebelle & Grudić, 2024). During the primary accretion phase (which includes most Class 0 and many Class I objects), protostars amass the bulk of their material due to high accretion rates (e.g., Watson et al., 2016; Manoj et al., 2016; Fischer et al., 2017; Federman et al., 2023; Narang et al., 2023). It is also during the primary accretion phase that the protostellar disk is formed, and the initial conditions of planet formation are set (Tobin & Sheehan, 2024).
@@
Protostars, regardless of their mass (spanning three orders of magnitude), share a common structural blueprint (Beltrán & de Wit, 2016; Tobin & Sheehan, 2024), featuring a central protostar surrounded by a dense infalling envelope of gas and dust, and a circumstellar disk. Collimated jets and wide-angle winds are also launched from the disk that carves cavities into the envelope, launching outflows of entrained gas and reducing the mass available to the nascent protostar (Lee, 2020; Ray & Ferreira, 2021; Habel et al., 2021; Hsieh et al., 2023; Shang et al., 2023b; Dunham et al., 2024). At the same time these jets and winds remove angular momentum from the disk, thereby driving accretion onto the central protostar (e.g., Shang et al., 2007; Hartmann et al., 2016; Pudritz & Ray, 2019; Ray & Ferreira, 2021; Tabone et al., 2022).
@@
Among the various proposed mechanisms for launching outflows, disk winds and X-winds are considered to be the leading contenders. Disk wind models have a characteristic nested velocity structure, where more collimated flows exhibit higher velocities (e.g., Pudritz & Norman, 1983, 1986; Shu et al., 1994; Cabrit et al., 1999; Shang et al., 2007; de Valon et al., 2022; Pascucci et al., 2023; Oliva & Kuiper, 2023; Feeney-Johansson et al., 2025). JWST observations of protostars and Class II disks also appear to find such nested outflows in atomic and H2 emission from YSOs (e.g., Pascucci et al., 2025; Caratti o Garatti et al., 2024; Delabrosse et al., 2024; Tychoniec et al., 2024, Narang et al., in prep, Tyagi et al., in prep); but also see Narang et al. (2025). In contrast, X-winds have a similar velocity for the collimated and wide (opening) angle component, although recent models of X-winds generate a slower, wide angle component of shock compressed wind (Shang et al., 2020, 2023a). Recently Nakatani et al. (2025) have demonstrated that photoevaporative wind models can also reproduce several key properties of the observed H2 winds, such as their nested morphology and overall flux levels; however, these models generally underpredict the high- emission associated with more spatially extended H2 structures.
@@
A distinctive feature of disk winds is that they originate across a broad range of disk radii, from the inner disk regions around au to several 10s of au (Pudritz & Norman, 1983, 1986; Wardle & Koenigl, 1993; Konigl & Pudritz, 2000; Pudritz et al., 2007; Wang et al., 2019; Pascucci et al., 2023; Oliva & Kuiper, 2023), with velocities ranging from a few km s-1 to several tens of km s-1. In contrast, X-winds are launched when stellar magnetic field lines connect with the disk at the co-rotation radius (around au), generating fast ( km s-1) MHD winds that spread widely (e.g., Shu et al., 1994; Shang et al., 2007; Pascucci et al., 2023). Another class of models to explain the outflows from protostars is the jet-driven bow-shock (e.g., Masson & Chernin, 1993; Raga & Cabrit, 1993; Ostriker et al., 2001; Rabenanahary et al., 2022). In this scenario, the jet propagates into the surrounding cloud and forms bow shocks that accelerate and push the ambient gas, producing outflow shells surrounding the jet. Models invoking a pulsed jet create internal working surfaces in the jets that may entrain and accelerate gas inside the cavity, which may appear similar to a disk wind (Rabenanahary et al., 2022; Rivera-Ortiz et al., 2023).
@@ IV Origin of the outflow from IRAS 16253-2429
The MHD disk wind naturally explains the presence of a primarily molecular component of the flow (Panoglou et al., 2012). Evidence for disk winds being the driving mechanism behind the nested structure of IRAS 16253-2429 comes from the analysis of the kinematics of the molecular outflow as traced in H2. The position-velocity diagrams reveal a clear distinction in the kinematics of the low- and high-excitation H2 transitions. The low- H2 transitions exhibit relatively low velocities, 12.5 km s-1 with respect to the mean flow velocity. In contrast, the higher- transitions, such as H2 0-0 S(5) and S(7), display significantly higher velocities, around 28.5 km s-1 with respect to the mean flow velocity. This suggests that the higher-energy H2 emission, not only appears more collimated but is also moving considerably faster compared to the wider, slower-moving low- H2 emission. This is consistent with the results of MHD wind simulations, which predicts faster speeds for more collimated parts of the flow (e.g.,Pudritz & Norman 1983, 1986; Wardle & Koenigl 1993; Konigl & Pudritz 2000; Pudritz et al. 2007; Pascucci et al. 2023).
In contrast, X-wind models predict that both the jet and the wide-angle wind components are launched from the innermost regions of the disk and therefore should exhibit similar velocities. Within the X-wind framework (Najita & Shu, 1994; Shang et al., 2007), the slower H2 emission we observe would be tracing shocked material within the wind. To reproduce the observed flow, X-wind models must account for both the molecular nature of the warm, slower component and the atomic or ionic composition of the fast, collimated jet, which some recent simulations have begun to achieve (e.g., Shang et al., 2020, 2023a) by introducing the interaction of X-winds with the ambient material.
Recently, Nakatani et al. (2025) demonstrated that photoevaporative winds can exhibit morphological characteristics similar to those of MHD winds, with higher- H2 transitions displaying a more collimated structure. However, MHD disk–wind models predict predominantly molecular outflows (e.g., Panoglou et al., 2012), whereas the strong high-energy radiation required to drive photoevaporative winds can efficiently dissociate H2, resulting in very low molecular fractions (Wang & Goodman, 2017). IRAS 16253-2429 is currently in a quiescent phase and is therefore unlikely to possess a sufficiently strong UV field to drive photoevaporation. Moreover, the models of Nakatani et al. (2025) predict that photoevaporative winds typically reach velocities of , whereas for IRAS 16253-2429 we measure significantly higher velocities, up to , albeit with large uncertainties. Finally, the presence of highly collimated atomic/ionic jets traced by e.g., [Fe II], [Ne II], and [Ar II] (Narang et al., 2024) further supports an MHD wind scenario over a photoevaporative origin.

### v2

@@ I Introduction
The process of star formation is governed by the complex interplay between infall, accretion, and feedback mechanisms, which determine the final mass of the star and the initial conditions for planet formation (Dunham et al., 2014; Federrath et al., 2014; Guszejnov et al., 2018, 2022; Fischer et al., 2023; Hsieh et al., 2023; Tobin and Sheehan, 2024; Hennebelle and Grudić, 2024). During the primary accretion phase (which includes most Class 0 and many Class I objects), protostars amass the bulk of their material due to high accretion rates (e.g., Watson et al., 2016; Manoj et al., 2016; Fischer et al., 2017; Federman et al., 2023; Narang et al., 2023). It is also during the primary accretion phase that the protostellar disk is formed, and the initial conditions of planet formation are set (Tobin and Sheehan, 2024).
@@
Protostars, regardless of their mass (spanning three orders of magnitude), share a common structural blueprint (Beltrán and de Wit, 2016; Tobin and Sheehan, 2024), featuring a central protostar surrounded by a dense infalling envelope of gas and dust, and a circumstellar disk. Collimated jets and wide-angle winds are also launched from the disk that carves cavities into the envelope, launching outflows of entrained gas and reducing the mass available to the nascent protostar (Lee, 2020; Ray and Ferreira, 2021; Habel et al., 2021; Hsieh et al., 2023; Shang et al., 2023b; Dunham et al., 2024). At the same time these jets and winds remove angular momentum from the disk, thereby driving accretion onto the central protostar (e.g., Shang et al., 2007; Hartmann et al., 2016; Pudritz and Ray, 2019; Ray and Ferreira, 2021; Tabone et al., 2022).
@@
Among the various proposed mechanisms for launching outflows, disk winds and X-winds are considered to be the leading contenders. Disk wind models have a characteristic nested velocity structure, where more collimated flows exhibit higher velocities (e.g., Pudritz and Norman, 1983, 1986; Shu et al., 1994; Cabrit et al., 1999; Shang et al., 2007; de Valon et al., 2022; Pascucci et al., 2023; Oliva and Kuiper, 2023; Feeney-Johansson et al., 2025). JWST observations of protostars and Class II disks also appear to find such nested outflows in atomic and H2 emission from YSOs (e.g., Pascucci et al., 2025; Caratti o Garatti et al., 2024; Delabrosse et al., 2024; Tychoniec et al., 2024, Narang et al., in prep, Tyagi et al., in prep); but also see Narang et al. (2025). In contrast, X-winds have a similar velocity for the collimated and wide (opening) angle component, although recent models of X-winds generate a slower, wide angle component of shock compressed wind (Shang et al., 2020, 2023a). Recently Nakatani et al. (2025) have demonstrated that photoevaporative wind models can also reproduce several key properties of the observed H2 winds, such as their nested morphology and overall flux levels; however, these models generally underpredict the high- emission associated with more spatially extended H2 structures.
@@
A distinctive feature of disk winds is that they originate across a broad range of disk radii, from the inner disk regions around au to several 10s of au (Pudritz and Norman, 1983, 1986; Wardle and Koenigl, 1993; Konigl and Pudritz, 2000; Pudritz et al., 2007; Wang et al., 2019; Pascucci et al., 2023; Oliva and Kuiper, 2023), with velocities ranging from a few km s-1 to several tens of km s-1. In contrast, X-winds are launched when stellar magnetic field lines connect with the disk at the co-rotation radius (around au), generating fast ( km s-1) MHD winds that spread widely (e.g., Shu et al., 1994; Shang et al., 2007; Pascucci et al., 2023). Another class of models to explain the outflows from protostars is the jet-driven bow-shock (e.g., Masson and Chernin, 1993; Raga and Cabrit, 1993; Ostriker et al., 2001; Rabenanahary et al., 2022). In this scenario, the jet propagates into the surrounding cloud and forms bow shocks that accelerate and push the ambient gas, producing outflow shells surrounding the jet. Models invoking a pulsed jet create internal working surfaces in the jets that may entrain and accelerate gas inside the cavity, which may appear similar to a disk wind (Rabenanahary et al., 2022; Rivera-Ortiz et al., 2023).
@@ IV Origin of the outflow from IRAS 16253-2429
The MHD disk wind naturally explains the presence of a primarily molecular component of the flow (Panoglou et al., 2012). Evidence for disk winds being the driving mechanism behind the nested structure of IRAS 16253-2429 comes from the analysis of the kinematics of the molecular outflow as traced in H2. The position-velocity diagrams reveal a clear distinction in the kinematics of the low- and high-excitation H2 transitions. The low- H2 transitions exhibit relatively low velocities, 12.5 km s-1 with respect to the mean flow velocity. In contrast, the higher- transitions, such as H2 0-0 S(5) and S(7), display significantly higher velocities, around 28.5 km s-1 with respect to the mean flow velocity. This suggests that the higher-energy H2 emission, not only appears more collimated but is also moving considerably faster compared to the wider, slower-moving low- H2 emission. This is consistent with the results of MHD wind simulations, which predicts faster speeds for more collimated parts of the flow (e.g.,Pudritz and Norman 1983, 1986; Wardle and Koenigl 1993; Konigl and Pudritz 2000; Pudritz et al. 2007; Pascucci et al. 2023).
In contrast, X-wind models predict that both the jet and the wide-angle wind components are launched from the innermost regions of the disk and therefore should exhibit similar velocities. Within the X-wind framework (Najita and Shu, 1994; Shang et al., 2007), the slower H2 emission we observe would be tracing shocked material within the wind. To reproduce the observed flow, X-wind models must account for both the molecular nature of the warm, slower component and the atomic or ionic composition of the fast, collimated jet, which some recent simulations have begun to achieve (e.g., Shang et al., 2020, 2023a) by introducing the interaction of X-winds with the ambient material.
Recently, Nakatani et al. (2025) demonstrated that photoevaporative winds can exhibit morphological characteristics similar to those of MHD winds, with higher- H2 transitions displaying a more collimated structure. However, MHD disk–wind models predict predominantly molecular outflows (e.g., Panoglou et al., 2012), whereas the strong high-energy radiation required to drive photoevaporative winds can efficiently dissociate H2, resulting in very low molecular fractions (Wang and Goodman, 2017). IRAS 16253-2429 is currently in a quiescent phase and is therefore unlikely to possess a sufficiently strong UV field to drive photoevaporation. Moreover, the models of Nakatani et al. (2025) predict that photoevaporative winds typically reach velocities of , whereas for IRAS 16253-2429 we measure significantly higher velocities, up to , albeit with large uncertainties. Finally, the presence of highly collimated atomic/ionic jets traced by e.g., [Fe II], [Ne II], and [Ar II] (Narang et al., 2024) further supports an MHD wind scenario over a photoevaporative origin.